OpenStudy (anonymous):

I have 180 meters of string, I need to make a maximum area rectangle shape out of it. How the formula for area should sound like?

Vocaloid (vocaloid):

perimeter = 2L + 2W with me so far?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

Is this calculus ?

Vocaloid (vocaloid):

then if 180 = 2W + 2L then 90 = W + L, with me so far?

OpenStudy (anonymous):

Yes, then area= x(x-90) right?

Vocaloid (vocaloid):

almost

Vocaloid (vocaloid):

90 = W + L means that W = 90 - L so area = (90-L)*L (or x in your case)

OpenStudy (anonymous):

so then S(x)=-x^2 +90x, and the maximum I get is 90 which for me doesn't make sense

OpenStudy (phi):

you should take the derivative of your function, and set it equal to zero or, if you notice y= -x^2+90x has the shape of a parabola, you are finding the peak (i.e. vertex of this \(\cap\) shape) where the slope is 0